A dipole consists of 2 charges +10 μC and -10μC separated by a certain distance.Let they be located at x=6.0 cm,y=0 and x=-6.0 cm and y=0 respectively.Calculate the field strength at point x=0,y=8 cm.

 

OPTIONS:

  • 1.08×10⁵Nc⁻¹

  • 2.16×10⁵Nc⁻¹

  • 1.08×10⁷Nc⁻¹

  • 0.54×10⁵Nc⁻¹

  • 1.54×10⁵Nc⁻¹

Grade:12th Pass

1 Answers

Arun
5 years ago
Dear Ashwin
 
Electric field at the centre will be E = 2* kq/r² = 0.5 * 10^7 V/m.
Along y axis:
x = sqrt(36 +64) = 10 cm
 
E due to positive = E due to negative = kq/x²
 
E(net) = 2E cos theta = 2 (kq/x²) * (r/x)
 
 = 108 * 10^5 = 1.08 * 10^7 Volt/m